THE ROTATIONALLY RESOLVED ELECTRONIC SPECTRUM OF THE DIFLUOROBENZENE-N_{2} VAN DER WAALS COMPLEX^{a}

\underline{CHEOLHWA KANG} AND DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

The rotationally resolved S_{1}←S_{0} fluorescence excitation spectra of para-difluorobenzene (pDFB) and its N_{2} complex have been observed. The high resolution spectrum of pDFB is pure b-type and spans an unusually wide energy region of over 20 cm^{−1}. The N_{2} complex of pDFB exhibits two absorption bands shifted to the red of the bare molecule origin, at -27 and -25cm^{−1}. Each of these bands is a superposition of the two pure c-type spectra that result from the internal rotation of the N_{2} molecule in the complex. Analyses of these bands show that the N_{2} molecule lies ˜ 3.59Å above the center of the ring plane in the S_{0} state; this distance decreases to ˜ 3.47Å in the S_{1} state. The barriers to internal rotation in the two electronic states are also different. A full discussion of the factors responsible for the differences will be given.

awork supported by NSF